# Phantom Crossing and Oscillating Dark Energy with $F(R)$ Gravity

### v1

@@ IV Several Models that Exhibit Phantom Crossing: A Theoretical Account
Here $\Lambda_{0}$ and $\Lambda_{1}$ are constants with the dimension of $[m]^{4}$, and $C$ is a dimensionless constant. The effective energy density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ has a peak at $N=N_{0}$, which corresponds to the $e$-foldings that the inverse phantom crossing occurs.
@@ V Oscillating Hubble Rate
As in Elizalde:2011ds ; Elizalde:2010ts , we consider the model where the Hubble rate $H$ oscillates and therefore the Universe iterates the phantom crossing and the inverse phantom crossing. In this section, we consider such an oscillation for the total energy density or the behaviour of $H$. A simple model of this sort is given by,
@@
where $H_{0}$, $C$, and $\omega$ are constants, and we assume $10<C<$ 1 and $\omega>0$. Then we find,
@@
Because we are considering the expanding Universe, where $H>0$, when $-\frac{\pi}{2}+2n\pi<\omega N<\frac{\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is positive and therefore, the Universe is in the phantom phase. On the other hand, when $\frac{\pi}{2}+2n\pi<\omega N<\frac{3\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is negative and therefore, the Universe is in the non-phantom phase. The phantom crossing occurs when $\omega N=-\frac{\pi}{2}+2n\pi$ with an integer $n$ and the inverse phantom crossing occurs when $\omega N=\frac{\pi}{2}+2n\pi$ with an integer $n$. Under the assumption (58), the scalar curvature $R$ is given by
$R=12H^{2}+6HH^{\prime}=3{H_{0}}^{2}\left(4+\sqrt{16+\omega^{2}}C\sin\left(\omega N+\theta_{0}\right)\right)\,.$ (60)
Here $\theta_{0}$ is defined by $\tan\theta_{0}=\frac{\omega}{4}$. Eq. (60) can be solved with respect to $N$,
@@
Here $sin^{-1}$ is the inverse function of $\sin$ function. The $e$-foldings $N$ is explicitly given as a function of $R$. Furthermore, we find,
$\sin\left(\omega N\right)=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\,,$
$\cos\left(\omega N\right)=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,,$ (62)
@@
$m(N)=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{3N}}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}\,,$
$f=$ $\,-R-2\kappa^{2}\Phi_{0}{H_{0}}^{2}\left[3+C\left\{\left(3+\frac{9\omega^{2}}{2n_{\mathrm{DM\,(0)}}}\right)\sin\left(\omega N\right)\right.\right.$
@@
$m(R)=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{-\frac{3\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)}}{n_{\mathrm{DM\,(0)}}\sqrt{16+\omega^{2}}}\left\{3\omega^{2}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right.$
@@
$f(R)=$ $\,-R-2\kappa^{2}\Phi_{0}{H_{0}}^{2}\left[3+\frac{C}{\sqrt{16+\omega^{2}}}\left\{\left(12+\frac{3}{2}\omega^{2}+\frac{3\omega^{2}}{n_{\mathrm{DM\,(0)}}}\right)\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right.\right.$
@@ VI Scalaron as Dark Matter with Curvature-dependent Mass
The action of the $F(R)$ gravity is given by replacing the scalar curvature $R$ in the Einstein-Hilbert action, which is,
@@
We obtain $A=R$ by the variation of the action with respect to $A$ and by substituting the obtained equation $A=R$ into the action (70), we find that the action in (66) is reproduced. If we rescale the metric by a kind of scale transformation,
@@
Therefore, when $R>4\Lambda+\frac{1}{2\alpha}$, the scalar mode becomes tachyonic. This is not surprising because it corresponds to the slow-roll during inflation. The model (74) could describe inflation due to the last term, $\alpha R^{2}$, when $R$ is large. The second term could play the role of a small cosmological term for the late-time accelerated expansion, when $R$ is small. Then we expect $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ and $\Lambda\sim\left(10^{-33}\,\mathrm{eV}\right)^{2}$. Therefore, we may estimate $m_{\sigma}$ as $m_{\sigma}\sim 10^{13}\,\mathrm{GeV}$, which is large enough. Anyway, the model (74) simply describes both the inflation and the DE.
@@ VII Towards a Realistic Model: A Theoretical Approach
In order to obtain a realistic model, we need to clarify the creation of the scalaron particles and to check the consistency of the parameters. First, we should note that, as is clear from the scale transformation (71), there is an interaction between the scalaron and the matter, including the standard model particles. Then, in the thermal equilibrium, the scalaron could be created by the interaction. Because the coupling is given by the gravitational coupling, which is the inverse of the Planck mass, the interaction is very weak, and therefore, a sufficient amount of the scalaron particles could be created, and the mass of the scalaron must be small enough. Eq. (75) could tell that the mass might be small just after the inflationary era. There might also be gravitational particle production.
We should note that the DESI observation corresponds to the redshift region $3.50<z<$ 3.5. If we choose the parameter $\alpha$ in the model in (74) to be $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ so that the model could also describe the inflation, Eq. (VI) tells that in the DESI redshift, the mass is already heavy enough, ${m_{\sigma}}^{2}\sim\frac{3}{4\alpha}$, and therefore the inverse phantom crossing could not occur. In order to avoid this problem, we may choose $1/\alpha$ to be small enough. In the laboratory, the energy density of the redshift region $3.50<z<$ 3.5 can be easily realised, and the fifth force experiment excludes such a light mode. The coupling between the scalaron and the matter shifts the mass due to the Chameleon mechanism, but the scalaron could be light enough to be excluded from the experiments. In order to avoid the problem, we may modify the model so that $\alpha$ could depend on the scalar curvature. The problem might be solved if $1/\alpha$ is large enough, even in the redshift region $3.50<z<$ 3.5 to be consistent with the fifth force experiment, and after that $1/\alpha$ further increases to generate the apparent phantom crossing.
@@ VIII Some Explicit Examples of Oscillating DE with Phantom Crossing
Now let us consider some realistic models of $F(R)$ gravity DE which we will extensively study using numerical methods, and we will obtain the cosmological parameters of these models. We will be mainly interested in the DE density and DE EoS parameter, which are constrained by the data, and we will numerically show that the models exhibit late-time EoS oscillations and a phantom crossing. Let us consider the $F(R)$ gravity action in the presence of perfect matter fluids in (66). Again we assume that the $F(R)$ gravity has the form of (4) and upon varying the action (66) with respect to the FLRW metric, we obtain,
@@
where $\rho_{\mathrm{total}}$ stands for the total energy density of the total effective cosmological fluid, which includes the $F(R)$ gravity contribution, and $p_{\mathrm{total}}$ denotes the total pressure. The total cosmological fluid is composed of three parts: the cold dark matter one ($\rho_{\mathrm{matter}}$), the radiation ($\rho_{\mathrm{r}}$) and the geometric part of $F(R)$ gravity ($\rho_{\mathrm{DE}}$). Therefore, we have, $\rho_{\mathrm{total}}=\rho_{\mathrm{matter}}+\rho_{\mathrm{r}}+\rho_{\mathrm{DE}}$ and in addition $p_{\mathrm{total}}=p_{\mathrm{matter}}+p_{\mathrm{r}}+p_{\mathrm{DE}}$. It is the geometric fluid which drives the dynamical evolution in the late-time era, and with its energy density and pressure being,
@@
and in addition, we introduce the statefinder parameter $y_{H}(z)$ Hu:2007nk ; Bamba:2012qi ; Nojiri:2017ncd ,
@@
$J_{1}=$ $\,\frac{1}{(z+1)}\left(-3-\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{1-F_{R}}{6{m_{s}}^{2}F_{RR}}\right)\,,$ (86)
$J_{2}=$ $\,\frac{1}{(z+1)^{2}}\left(\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{2-F_{R}}{3{m_{s}}^{2}F_{RR}}\right)\,,$ (87)
$J_{3}=$ $\,-3(z+1)-\frac{(1-F_{R})\left((z+1)^{3}+2\chi(z+1)^{4}\right)+(R-F)/\left(3{m_{s}}^{2}\right)}{(z+1)^{2}\left(y_{H}+(z+1)^{3}+\chi(z+1)^{4}\right)}\frac{1}{6{m_{s}}^{2}F_{RR}}\,,$ (88)
@@
$y_{H}(z_{f})=\frac{\Lambda}{3{m_{s}}^{2}}\left(1+\frac{1+z_{f}}{1000}\right)\,,\quad\left.\frac{dy_{H}(z)}{dz}\right|_{z=z_{f}}=\frac{1}{1000}\frac{\Lambda}{3{m_{s}}^{2}}\,,$ (91)
@@
where the “prime” denotes differentiation with respect to the redshift. In addition, and for comparison reasons, let us quote here the Hubble rate for the $\Lambda$CDM model,
@@
where $\Omega_{\Lambda}\simeq 0.68136$ and $\Omega_{\mathrm{matter}}\simeq 0.3153$ and $H_{0}\simeq 1.37187\times 10^{-33}\,\mathrm{eV}$ according to the latest 2018 Planck data Planck:2018vyg . Now let us focus on the phenomenology of two interesting models which exhibit remarkable behaviour at late times. We start off with the following model,
@@
with $b$ and $R_{0}$ are free parameters with mass dimensions $[m]^{2}$ and also $c$ is a dimensionless parameter. We obtained a viable phenomenology for the choice $b=20\Lambda,$ $c=2$, and $R_{0}={m_{s}}^{2}/0.00091$. In Fig. 1 we plot the deceleration parameter for this model (red curve) compared with the $\Lambda$CDM (black curve), and the DE EoS for various redshift ranges. As can be seen, the DE EoS is strongly oscillating at late times, and the model ends up in a quintessential regime. The DE energy density parameter and the DE EoS parameter at $z=0$ are for this model $\Omega_{\mathrm{DE}}=0.6918$ and $w_{\mathrm{DE}}=-0.9974$ which comply with the latest Planck constraints.
@@
with $b$, $c$, $d$, $n$ being dimensionless parameters and $\alpha$, $R_{0}$ are also free parameters with mass dimensions $[m]^{2}$. With the choice $\alpha=1.4\Lambda$, $b=1$, $c=0.2$, $d=0.04$, $R_{0}={m_{s}}^{2}$, and $n=0.3$, we obtain an interesting DE oscillating behaviour which is also phenomenologically viable. Specifically, in Fig. 2, we plot the deceleration parameter (red curve) compared with the $\Lambda$CDM (black curve), and also the DE EoS for various redshift ranges, focusing on small redshifts. As can be seen in this case, too, the DE EoS is strongly oscillating at late times and at small redshifts before the present day, and the model ends up in a quintessential regime in this case, too. The DE energy density parameter and the DE EoS parameter at $z=0$ for this model are $\Omega_{\mathrm{DE}}=0.6851$ and $w_{\mathrm{DE}}=-0.9887$ which are again compatible with the latest Planck constraints.
@@
Thus in this section, using a numerical approach, we demonstrated that it is possible to obtain oscillating DE eras using rather simple $F(R)$ gravity models. Note that the models presented in this section were also extensively examined regarding their phenomenology in Ref. Odintsov:2025jfq and the model (99) is found to be fully compatible with the observational data, including the latest DESI data.

### v2

@@ IV Several Models that Exhibit Phantom Crossing: A Theoretical Account
Here $\Lambda_{0}$ and $\Lambda_{1}$ are constants with the dimension of $m]^{4}$, and $C$ is a dimensionless constant. The effective energy density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ has a peak at $N=N_{0}$, which corresponds to the $e$-foldings that the inverse phantom crossing occurs.
@@ V Oscillating Hubble Rate
As in Elizalde:2011ds ; Elizalde:2010ts , we consider the model where the Hubble rate $H$ oscillates and therefore the Universe iterates the phantom crossing and the inverse phantom crossing. In this section, we consider such a oscillation for the total energy density or the behavior of $H$. A simple model of this sort is given by,
@@
where $H_{0}$, $C$, and $\omega$ are constants, and we assume $0<C<1$ and $\omega>0$. Then we find,
@@
Because we are considering the expanding Universe, where $H>0$, when $\frac{\pi}{2}+2n\pi<\omega N<\frac{\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is positive and therefore, the Universe is in the phantom phase. On the other hand, when $\frac{\pi}{2}+2n\pi<\omega N<\frac{3\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is negative and therefore, the Universe is in the non-phantom phase. The phantom crossing occurs when $\omega N=-\frac{\pi}{2}+2n\pi$ with an integer $n$ and the inverse phantom crossing occurs when $\omega N=\frac{\pi}{2}+2n\pi$ with an integer $n$. Under the assumption (58), the scalar curvature $R$ is given by
$R=$  $\,12H^{2}+6HH^{\prime}=H_{0}^{2}\left(12+12C\sin\left(\omega N\right)+3C\omega\cos\left(\omega N\right)\right)$
$=$  $\,3{H_{0}}^{2}\left\{4+\sqrt{16+\omega^{2}}C\left(\frac{4}{\sqrt{16+\omega^{2}}}\sin\left(\omega N\right)+\frac{\omega}{\sqrt{16+\omega^{2}}}\cos\left(\omega N\right)\right)\right\}$
$=$  $\,3{H_{0}}^{2}\left(4+\sqrt{16+\omega^{2}}C\sin\left(\omega N+\theta_{0}\right)\right)\,.$ (60)
Here $\theta_{0}$ is defined by $\tan\theta_{0}=\frac{\omega}{4}$. Eq. (V) can be solved with respect to $N$,
@@
Here $sin^{-1}$ is the inverse function of $\sin$ function. The $e$-foldings NN is explicitly given as a function of $R$. Furthermore, we find,
$\sin\left(\omega N\right)=$ $\,\sin\left(\omega N+\theta_{0}-\theta_{0}\right)=\sin\left(\omega N+\theta_{0}\right)\cos\left(\theta_{0}\right)-\cos\left(\omega N+\theta_{0}\right)\sin\left(\theta_{0}\right)$
$=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\,$,
$\cos\left(\omega N\right)=$ $\,\cos\left(\omega N+\theta_{0}-\theta_{0}\right)=\cos\left(\omega N+\theta_{0}\right)\cos\left(\theta_{0}\right)+\sin\left(\omega N+\theta_{0}\right)\sin\left(\theta_{0}\right)$
$=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,$, (62)
@@
$m(N)=$ $\,\frac{\Phi_{0}{H_{0}}^{2}\mathrm{e}^{3N}}{3n_{\mathrm{DM\,(0)}}}\left\{-3\left(C\omega\cos\left(\omega N\right)-\frac{C\omega^{2}}{2}\sin\left(\omega N\right)\right)-12C\omega\cos\left(\omega N\right)+3C\omega^{2}\sin\left(\omega N\right)\right\}$
$=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{3N}}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}\,$,
$f=$ $\,-R+2\kappa^{2}\left\{-3\Phi_{0}{H_{0}}^{2}\left(1+C\sin\left(\omega N\right)+\frac{C\omega}{2}\cos\left(\omega N\right)\right)\right\}$
$\,-\frac{2\kappa^{2}\Phi_{0}{H_{0}}^{2}C}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}$
[…]
@@
[…]
$\,\left.-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}-15\omega\left\{\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right$.
$\,\left.\left.+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right\}\right]\,$,
$=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{-\frac{3\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)}}{n_{\mathrm{DM\,(0)}}\sqrt{16+\omega^{2}}}\left\{3\omega^{2}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right$.
@@
[…]
$\,\left.-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}$
$\,+\left(\frac{3}{2}-\frac{15}{n_{\mathrm{DM\,(0)}}}\right)\omega\left\{\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right$.
$\,\left.\left.+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right\}\right]$
[…]
@@ VI Scalaron as Dark Matter with Curvature-dependent Mass
The action of the $F(R)$ gravity is given by replacing the scalar curvature $R$ in the Einstein-Hilbert action which is,
@@
We obtain $A=R$ by the variation of the action with respect to $A$ and by substituting the obtained equation $A=R$ into the action (70), we find that the action in (66) is reproduced. If we rescale the metric by a kind of a scale transformation,
@@
Therefore, when $R>4\Lambda+\frac{1}{2\alpha}$, the scalar mode becomes tachyonic. This is not surprising because it corresponds to the slow-roll during inflation. The model (74) could describe inflation due to the last term, $\alpha R^{2}$, when $R$ is large. The second term could play the role of a small cosmological term for the late-time accelerated expansion, when $R$ is small. Then we expect $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ and $\Lambda\sim\left(10^{-33}\,\mathrm{eV}\right)^{2}$. Therefore, we may estimate $m_{\sigma}$ as $m_{\sigma}\sim 10^{13}\,\mathrm{GeV}$, which is large enough. In any case, the model (74) simply describes both the inflation and the DE.
@@ VII Towards a Realistic Model: A Theoretical Approach
In order to obtain a realistic model, we need to clarify the creation of the scalaron particles and to check the consistency of the parameters. First, we should note that, as is clear from the scale transformation (71), there is an interaction between the scalaron and the matter, including the standard model particles. Then, in the thermal equilibrium, the scalaron could be created by the interaction. Because the coupling is given by the gravitational coupling, which is the inverse of the Planck mass, the interaction is very weak and therefore a sufficient amount of the scalaron particles could be created, and the mass of the scalaron must be small enough. Eq. (75) could tell that the mass might be small just after the inflationary era. There might also be gravitational particle production.
We should note that the DESI observation corresponds to the redshift region $0<z<3.5$. If we choose the parameter $\alpha$ in the model in (74) to be $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ so that the model could also describe the inflationary era, Eq. (VI) tells that in the DESI redshift, the mass is already heavy enough, ${m_{\sigma}}^{2}\sim\frac{3}{4\alpha}$, and therefore the inverse phantom crossing could not occur. In order to avoid this problem, we may choose $1/\alpha$ to be small enough. In the laboratory, the energy density of the redshift region $0<z<3.5$ can be easily realized, and the fifth force experiment excludes such a light mode. The coupling between the scalaron and the matter shifts the mass due to the Chameleon mechanism, but the scalaron could be light enough to be excluded from the experiments. In order to avoid this problem, we may modify the model so that $\alpha$ could depend on the scalar curvature. The problem might be solved if $1/\alpha$ is large enough even in the redshift region $0<z<3.5$ to be consistent with the fifth force experiment and after that $1/\alpha$ further increases to generate the apparent phantom crossing.
@@ VIII Some Explicit Examples of Oscillating DE with Phantom Crossing
Now let us consider some realistic models of $F(R)$ gravity DE which we will extensively study using numerical methods and we will obtain the cosmological parameters of these models. We will be mainly interested in the DE density and DE EoS parameter which are constrained by the data and we will numerically show that the models exhibit late-time EoS oscillations and a phantom crossing. Let us consider the $F(R)$ gravity action in the presence of perfect matter fluids in (66). Again we assume that the $F(R)$ gravity has the form of (4) and upon varying the action (66) with respect to the FLRW metric, we obtain,
@@
where $\rho_{\mathrm{total}}$ stands for the total energy density of the total effective cosmological fluid, which includes the $F(R)$ gravity contribution, and $p_{\mathrm{total}}$ denotes the total pressure. The total cosmological fluid is composed by three parts, the cold dark matter one ($\rho_{\mathrm{matter}}$), the radiation ($\rho_{\mathrm{r}}$) and the geometric part of $F(R)$ gravity ($\rho_{\mathrm{DE}}$). Therefore, we have, $\rho_{\mathrm{total}}=\rho_{\mathrm{matter}}+\rho_{\mathrm{r}}+\rho_{\mathrm{DE}}$ and in addition $p_{\mathrm{total}}=p_{\mathrm{matter}}+p_{\mathrm{r}}+p_{\mathrm{DE}}$. It is the geometric fluid which drives the dynamical evolution in the late-time era, and with its energy density and pressure being,
@@
and in addition we introduce the statefinder parameter $y_{H}(z)$ Hu:2007nk ; Bamba:2012qi ; reviews1 ,
@@
$J_{1}=\frac{1}{(z+1)}\Big(-3-\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{1-F_{R}}{6{m_{s}}^{2}F_{RR}}\Big)\,$, (86)
$J_{2}=\frac{1}{(z+1)^{2}}\Big(\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{2-F_{R}}{3{m_{s}}^{2}F_{RR}}\Big)\,$, (87)
$J_{3}=-3(z+1)-\frac{(1-F_{R})((z+1)^{3}+2\chi(z+1)^{4})+(R-F)/(3{m_{s}}^{2})}{(z+1)^{2}(y_{H}+(z+1)^{3}+\chi(z+1)^{4})}\frac{1}{6{m_{s}}^{2}F_{RR}}\,$, (88)
@@
$y_{H}(z_{f})=\frac{\Lambda}{3{m_{s}}^{2}}\Big(1+\frac{1+z_{f}}{1000}\Big)\ ,\ \frac{dy_{H}(z)}{dz}\Big|_{z=z_{f}}=\frac{1}{1000}\frac{\Lambda}{3{m_{s}}^{2}},$ (91)
@@
[…]
@@
where $\Omega_{\Lambda}\simeq 0.68136$ and $\Omega_{\mathrm{matter}}\simeq 0.3153$ and $H_{0}\simeq 1.37187\times 10^{-33}\,\mathrm{eV}$ according to the latest 2018 Planck data Planck:2018vyg . Now let us focus on the phenomenology of two interesting models which exhibit remarkable behavior at late times. We start off with the following model,
@@
[…]
@@
[…]
@@
Thus in this section using a numerical approach, we demonstrated that it is possible to obtain oscillating DE eras using rather simple $F(R)$ gravity models. Note that the models presented in this section were also extensively examined regarding their phenomenology in Ref. Odintsov:2025jfq and the model (99) is found to be fully compatible with the observational data, including the latest DESI data.